Al's age plus Bob's age =33
Al's age plus Charlie's age =31
Bob's age plus Charlie's age = 26
How old is each boy?

Answer by Earlsdon(6294)

Let A = Al's age, B = Bob's age, and C = Charlie's age.
From the problem description you can write:
1) A+B = 33
2) A+C = 31
3) B+C = 26 Rewrite this as: C = 26-B and substitute into the 2nd equation and solve for A.
2a) A+(26-B) = 31
2b) A = 31-(26-B) Simplify.
2c) A = 5+B Now substitute this into the 1st equation and solve for B.
1a) (5+B)+B = 33
1b) 5+2B = 33 Subtract 5 from both sides.
1c) 2B = 28 Divide both sides by 2.
1d) B = 14 This is Bob's age. Substitute this into equation 2c) and solve for A.
2d) A = 5+14
2e) A = 19 This is Al's age. Substitute this into equation 2) and solve for C.
2f) 19+C = 31 Subtract 19 from both sides.
2g) C = 12 This is Charlie's age.